Check whether the sequence {an}, where


an= 1/(n+1)+1/(n+2)+..1/2n is convergent or not

1
Expert's answer
2022-03-01T12:12:20-0500

ANSWER. The sequence "\left \{ a_{n} \right \}" is convergent.

EXPLANATION.

Since "a_{n}=\sum_{k=1}^{n}\frac{1}{n+k}" , then "a_{n+1}=\sum_{k=1}^{n+1}\frac{1}{(n+1) +k}" and "a_{n+1} -a_{n}=\frac{1}{2n+1}+\frac{1}{2n+2}-\frac{1}{n+1}=\frac{1}{2n+1}-\frac{1}{2n+2} =\frac{1}{(2n+1)(2n+2)}>0" . Hence "a_{n+1}>a_{n }" . So , the sequence is increasing.

Because for all "k" such that "1\leq k\leq n" the inequality "(1+n)\leq(k+n)" is true and "\frac{1}{k+n}\leq\frac{1}{1+n}" . Therefore, "0<a_{n}\leq \sum_{k=1}^{n}\frac{1}{n+1}==\underset { n }{ \underbrace { \frac { 1 }{ n+1 } +...+\frac { 1 }{ n+1 } } }= \frac{n}{n+1} <1" .

Thus, it is proved that the increasing sequence is bounded from above, hence the sequence converges.